OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = 1/2x - 8. y =1/2x + 7 y = 1/2x + 5 y = 2x + 7 y = 2x + 5

OpenStudy (anonymous):

Parallel lines have the gradient. so Once its parallel to y = 1/2x - 8 then the gradient(m) = 1/2 Therefore using the formula for finding an equation of a line, y - y1 = m(x - x1), where x1 = -2 and y1 = 6 y -6 = (1/2)(x -(-2)) y -6 = (1/2)(x +2) y - 6 = 1/2 x +1 y = 1/2x +1 +6 Hence, y = 1/2x + 7 So answer is A.